What is the equation of a line that passes through (-2,-1) and (-4, -3)
Found 2 solutions by lwsshak3, stanbon:
Answer by lwsshak3(11628)

What is the equation of a line that passes through (-2,-1) and (-4, -3)
***
slope intercept form of equation for a line: y=mx+b, m=slope, b=y-intercept
For given problem:
m=∆y/∆x=(-3-(-1))/(-4-(-2)=-2/-2=1
equation: y=1x+b
using coordinates (-2, -1) on the line
b=y-x=-1+2=1
equation: y=x+1

Answer by stanbon(75887)

What is the equation of a line that passes through (-2,-1) and (-4, -3)
------------
slope = (-3--1)/(-4--2) = -2/-2 = 1
------------
Form:: y = mx + b
---
Solve for "b":
-1 = 1(-2) + b
b = 1
-------
Equation:
y = x + 1
-------------
Cheers,
Stan H.
-----------